OpenStudy (anonymous):

What is the exact value for the expression the square root of 153 minus the square root of 17 plus the square root of 68.

OpenStudy (anonymous):

@precal @jdoe0001

OpenStudy (precal):

try using the equation on the bottom to post your question, give it a try

OpenStudy (anonymous):

k

OpenStudy (anonymous):

\[\sqrt{153}-\sqrt{17}+\sqrt{68}\]

OpenStudy (anonymous):

do I do the factor tree @precal

OpenStudy (precal):

yes

OpenStudy (precal):

you are trying to find the same number under the radical

OpenStudy (anonymous):

Oh ok so 153 is divisble by 9

OpenStudy (precal):

yes and the square root of 9 is 3 so that would be a good choice

OpenStudy (anonymous):

9 and 17

OpenStudy (anonymous):

yup

OpenStudy (precal):

do the same to the rest and tell me your answer and I will double check you

OpenStudy (anonymous):

Hold on

OpenStudy (jdoe0001):

\(\bf { \sqrt{153}-\sqrt{17}+\sqrt{68} \\ \quad \\ {\color{brown}{ 153\to 3\cdot 3\cdot 17\to 3^2\cdot 17\qquad 68\to 2\cdot 2\cdot 17\to 2^2\cdot 17}}\quad thus \\ \quad \\ \sqrt{{\color{brown}{ 3^2\cdot 17}}}-\sqrt{17}+\sqrt{{\color{brown}{ 2^2\cdot 17}}}\implies 3\sqrt{17}-\sqrt{17}+2\sqrt{17}\implies ? }\)

OpenStudy (anonymous):

not sure if thats right... @jdoe0001

OpenStudy (jdoe0001):

hmmm well... what's.... wrong you think?

OpenStudy (precal):

yes doing good so far

OpenStudy (anonymous):

I have\[\sqrt[4]{17}\] and

OpenStudy (anonymous):

\[\sqrt{17}\]

OpenStudy (precal):

9 can be written as 3^2 square root of 9 is 3

OpenStudy (anonymous):

those re the closest

OpenStudy (precal):

3-1+2=2+2=4 so 4 times square root of 17

OpenStudy (anonymous):

o ok thanks! I might have one more that is giving me problems

OpenStudy (precal):

you are looking to create like terms, in this case we use the square root of 17 as our like term, then you just combine the coefficients.

OpenStudy (precal):

no problem, close this one out and post a new thread

OpenStudy (anonymous):

ok its like the first one we did